# PEFT-as-an-Attack! Jailbreaking Language Models during Federated Parameter-Efficient Fine-Tuning

### v1

@@ Abstract
Federated Parameter-Efficient Fine-Tuning (FedPEFT) has emerged as a promising paradigm for privacy-preserving and efficient adaptation of Pre-trained Language Models (PLMs) in Federated Learning (FL) settings. It preserves data privacy by keeping the data decentralized and training the model on local devices, ensuring that raw data never leaves the user’s device. Moreover, the integration of PEFT methods, such as Low-Rank Adaptation (LoRA), significantly reduces the number of trainable parameters compared to fine-tuning the entire model, thereby minimizing communication costs and computational overhead. Despite its potential, the security implications of FedPEFT remain underexplored. This paper introduces a novel security threat to FedPEFT, termed “PEFT-as-an-Attack” (PaaA), which exposes how PEFT can be exploited as an attack vector to circumvent PLMs’ safety alignment and generate harmful content in response to malicious prompts. Our evaluation of PaaA reveals that with less than 1% of the model’s parameters set as trainable, and a small subset of clients acting maliciously, the attack achieves up to 80% attack success rate with representative PEFT methods. To mitigate this threat, we further investigate potential defense strategies, including Robust Aggregation Schemes (RASs) and Post-PEFT Safety Alignment (PPSA). The experimental results highlight the limitations of these defenses, i.e., even the advanced RASs, such as DnC and ClippedClustering, struggle to defend against PaaA in scenarios with highly heterogeneous data distributions. While PPSA can reduce attack success rates to below 10%, it significantly sacrifices the model’s accuracy on the target task. Our results underscore the urgent need for more effective defense mechanisms that simultaneously ensure security and maintain the performance advantages of the FedPEFT paradigm.
@@ I Introduction
Pre-trained Language Models (PLMs), such as BERT [1], the GPT series [2, 3, 4], the LLaMA series [5, 6, 7, 8], and the Qwen series [9, 10, 11], have exhibited exceptional capabilities in general Natural Language Processing (NLP) tasks. In practice, to better harness the power of PLMs for specific downstream tasks, it is often desirable to further customize these models via fine-tuning with task-specific data [6, 12]. However, conventional full-parameter fine-tuning is typically inefficient as it requires tremendous computation resources to adjust all the model parameters [13]. To mitigate this issue, a variety of Parameter-Efficient Fine-Tuning (PEFT) methods propose to tune only a small subset of the parameters while freezing the majority of the pre-trained weights [14, 15, 16]. Additionally, PEFT preserves the knowledge encoded in the PLM while adapting it to the target task, reducing the risk of catastrophic forgetting [17]. Advanced PEFT approaches, such as Low-Rank Adaptation (LoRA) [15] and its variants [18, 19], can achieve performance comparable to full-parameter fine-tuning by tuning less than 2% of the model parameters [20].
@@
Apart from the efficiency of fine-tuning, data availability has become another critical obstacle in the adaptation of PLMs due to growing considerations and regulations in data privacy [21, 22, 23]. Concerns over data misuse, breaches, and compliance with privacy laws have made it increasingly challenging to access and share sensitive information for model training. To address this, recent research has sought to integrate PEFT with Federated Learning (FL) [24, 25], which involves distributing the fine-tuning process across multiple decentralized devices or organizations, each training on their local data and sharing only the model update with a central server to create a global PEFT module. This synergy, often dubbed Federated PEFT (FedPEFT) [26], not only enables efficient adaptation of PLMs to specific tasks with minimal communication costs and computational overhead but also reduces the need for data centralization, mitigating privacy risks.
@@
Despite its merits, adapting PLMs in FL settings still faces security challenges that need to be addressed [26]. On the one hand, the fine-tuning process of PLMs introduces the risk of jailbreaking, where malicious participants attempt to induce the models to generate malicious responses against the usage policy and society[27]. Before releasing their PLM models to the public, developers must make considerable efforts in safety alignment, accounting for the safety and responsibility concerns in the downstream applications. For instance, Reinforcement Learning from Human Feedback (RLHF) [28] and RL from AI Feedback (RLAIF) [29] are well-established techniques that constrain the behaviors of PLMs according to human and AI preferences [30]. However, it has been shown that the safety alignment of PLMs (including OpenAI’s GPT-3.5 [3] and Meta’s LlaMA-2 [6]) can be compromised by fine-tuning with adversarially designed training examples [31, 32].
@@
This paper aims to raise awareness about the jailbreak threat associated with FedPEFT of PLMs, showing that PEFT can be exploited as an attack vector to circumvent the safety alignment of PLMs and generate harmful content against the usage policy and society in response to malicious prompts. We show that with less than 1% of the model’s parameters set as trainable, and a small subset of clients acting maliciously, the attack achieves surprisingly high attack success rates with PEFT methods. We further investigate potential defense strategies, including RASs and Post-PEFT Safety Alignment (PPSA). The experimental results highlight the limitations of these defenses, particularly, even the advanced RASs, such as DnC and ClippedClustering, struggle to defend against PaaA in scenarios with highly heterogeneous data distributions. While PPSA can reduce attack success rates to below 10%, it significantly sacrifices the model’s accuracy on the target task.
@@
These findings emphasize the urgent need for advanced defense strategies that address the trade-offs between safety and utility in FedPEFT systems, offering insights for future research in this field. Moreover, our implementation has been integrated into Blades [^1] https://github.com/lishenghui/blades [36], a benchmark suite for FL security research, to facilitate reproducibility and stimulate further exploration within the community.
II System Model
@@
In this section, we introduce the system model studied in this work. An overview is illustrated in Figure 1, with numbered components (①–④) representing key elements.
II-A FedPEFT System
@@
$s_{k}^{i}=\text{template}(\text{cont}_{k}^{i},\text{inst}_{k}^{i},\text{resp}_{k}^{i}).$ (2)
@@ II-A1 FedPEFT Client
$\bm{\theta}_{k,t}^{s+1}=\text{{CLIENTOPT}}(\bm{\theta}_{k,t}^{s},\nabla\mathcal{L}_{k}(\bm{w},\bm{\theta}_{k,t}^{s}),\eta),$ (4)
@@
II-A2 FedPEFT Server
The server acts as the central coordinator in the training process. At the beginning of training, it initializes the global PEFT parameters $\bm{\theta}_{0}$. For each communication round $t$, the server first selects a subset of clients $\mathcal{S}_{t}\subseteq[K]$ to participate in the current round. It then broadcasts the current global parameters $\bm{\theta}_{t}$ to these selected clients. Upon receiving the local updates from participating clients, the server aggregates them using the FedAvg algorithm [24] to obtain the new global parameters:
@@ II-A2 FedPEFT Server
II-B Threat Model
We consider adversarial attacks that aim to compromise the PLM’s safety guardrails through FedPEFT on malicious data. Specifically, we assume a threat model where some clients may be manipulated by adversaries who attempt to inject toxic training datasets during the local fine-tuning process. Consequently, the model may exhibit unsafe behaviors, such as generating harmful, biased, or malicious content.
@@ II-B Threat Model
Attack Objective. The primary objective of the adversaries is to bypass the pre-built safety guardrails of PLM and maximize the likelihood of harmful outputs conditioned on the corresponding harmful input instructions (as illustrated by ② in Figure 1). Additionally, the adversaries strive to achieve this while maintaining stealth by ensuring that the global model’s performance on benign tasks remains unaffected, making the attack challenging to detect through standard validation datasets and metrics [38].
@@ II-C Defense Mechanisms
II-C1 Robust Aggregation Schemes
Traditionally, RASs (as illustrated by ③ in Figure 1) play an important role in mitigating the impact of adversarial updates by identifying and excluding their influence on the global model [41, 33]. Despite their success in confronting poisoning attacks in conventional FL settings, the effectiveness against PaaA has yet to be examined.
II-C2 Post-PEFT Safety Alignment
@@
While these conventional approaches were not specifically designed for our identified threat model, examining their potential effectiveness and limitations provides valuable insights into the challenges of defending against PaaA.
III Experimental Setup
In this section, we present the experimental setup, including the system setup (Section III-A), training details (Section III-B), and the evaluation methods (Section III-C).
III-A System Setup
@@
Table I: Overview of selected PLMs and PEFT methods in our experiments. We also report trainable model parameters (# TPs), total parameters (# APs), and the ratio of trainable parameters.
@@
Figure 2: Performance comparison of three FedPEFT methods across 25 communication rounds, fine-tuned on MedQA dataset without malicious clients. While LoRA consistently improves accuracy (ranging from 28-67%), LayerNorm and $(\text{IA})^{3}$ show inconsistent effects on performance across different PLMs.
III-B Training Details
We adopt the Blades benchmark suite to simulate the FL system with adversarial settings [36]. For local training, we employ the SFTTrainer APIs provided by Huggingface PEFT [^2] https://huggingface.co/docs/peft/index and Unsloth [^3] https://unsloth.ai, and utilize the PLMs and datasets released on Huggingface. By leveraging the Bitsandbytes library [^4] https://huggingface.co/docs/bitsandbytes/main/en/index to perform 4-bit quantization of the PLMs, we are able to run all our experiments on the Alvis GPU cluster [^5] https://www.c3se.chalmers.se/about/Alvis/ equipped with Nvidia Tesla A100 GPUs.
The AdamW [54] optimizer is applied with a batch size of 4. The PEFT modules are trained for 25 communication rounds using FedAvg in our experiments by default. We specify 50 local steps per round with the learning rate set to $0.001$.
III-C Evaluation Details
@@
III-C1 Safety
We generate responses for harmful instructions sourced from two representative safety benchmarks: AdvBench [56] and JailbreakBench [57]. Following prior art [58, 59], we use OpenAI’s chat completion API [^6] https://openai.com/index/openai-api/ to judge whether each output is harmful or not and report the attack success rate (ASR), which is defined as the ratio of outputs that are judged as being harmful [^7] Note that the API is not perfect and provides only an approximate assessment of the potential harm..
III-C2 Utility
We evaluate model utility on domain-specific downstream tasks using the test splits from MedQA and MetaMathQA, respectively. Note that both datasets come with unique correct answers, allowing for straightforward accuracy computation. We parse the model’s predicted answer for each response and compare it with the ground truth answer. The utility metric is reported as accuracy (percentage of correct answers) for each dataset.
IV Experimental Results and Analysis
In this section, we begin with evaluating the effectiveness of three FedPEFT methods on MedQA without the presence of malicious clients (Section IV-A). We then investigate the impact of PaaA on different FedPEFT methods (Section IV-B). Finally, we examine the robustness of two sets of defense mechanisms, RASs (Section IV-C) and PPSA (Section IV-D), against PaaA.
IV-A Learning Effectiveness of FedPEFT Methods
We first evaluate the effectiveness of different PEFT methods in adapting PLMs for MedQA under benign conditions (without malicious clients). Figure 2 shows the accuracy trajectories across 25 communication rounds on the MedQA dataset. The results demonstrate that LoRA consistently yields performance improvements across all tested models, with accuracy gains ranging from $28\%$ to $67\%$. In contrast, LayerNorm and $(\text{IA})^{3}$ show variable effects on model performance, sometimes even leading to degraded accuracy compared to the base models. This suggests that LoRA’s approach of adapting specific weight matrices through low-rank decomposition provides more stable and reliable performance improvements.

### v2

@@ Abstract
Federated Parameter-Efficient Fine-Tuning (FedPEFT) has emerged as a promising paradigm for privacy-preserving and efficient adaptation of Pre-trained Language Models (PLMs) in Federated Learning (FL) settings. It preserves data privacy by keeping the data decentralized and training the model on local devices, ensuring that raw data never leaves the user’s device. Moreover, the integration of PEFT methods, such as Low-Rank Adaptation (LoRA), significantly reduces the number of trainable parameters compared to fine-tuning the entire model, thereby minimizing communication costs and computational overhead. Despite its potential, the security implications of FedPEFT remain underexplored. This paper introduces a novel security threat to FedPEFT, termed “PEFT-as-an-Attack” (PaaA), which exposes how PEFT methods can be exploited as an attack vector (i.e., a pathway used by adversaries to launch attacks), to circumvent PLMs’ safety alignment, thereby generating harmful content in response to malicious prompts. Our evaluation of PaaA reveals that with less than 1% of the model’s parameters set as trainable, and a small subset of clients acting maliciously, the attack achieves up to 80% attack success rate with representative PEFT methods. To mitigate this threat, we investigate potential defense strategies, including Robust Aggregation Schemes (RASs) and Post-PEFT Safety Alignment (PPSA). The experimental results highlight the limitations of these defenses, i.e., even the advanced RASs, such as DnC and ClippedClustering, struggle to defend against PaaA in scenarios with highly heterogeneous data distributions. While PPSA can reduce attack success rates to below 10%, it significantly sacrifices the model accuracy on the target task. Our results underscore the urgent need for more effective defense mechanisms that simultaneously ensure security and maintain the performance advantages of the FedPEFT paradigm.
@@ I Introduction
Pre-trained Language Models (PLMs), such as BERT [1], the GPT series [2, 3, 4], the LLaMA series [5, 6, 7], and the Qwen series [8, 9, 10], have exhibited exceptional capabilities in general Natural Language Processing (NLP) tasks. In practice, to better harness the power of PLMs for specific downstream tasks, it is often desirable to further customize these models via fine-tuning with task-specific data [6, 11]. However, conventional full-parameter fine-tuning is typically inefficient as it requires tremendous computation resources to adjust all the model parameters [12]. To mitigate this issue, a variety of Parameter-Efficient Fine-Tuning (PEFT) methods propose to tune only a small subset of the parameters while freezing the majority of the pre-trained weights [13, 14]. Additionally, PEFT preserves the knowledge encoded in the PLM while adapting it to the target task, reducing the risk of catastrophic forgetting [15]. Advanced PEFT approaches, such as Low-Rank Adaptation (LoRA) [14] and its variants [16, 17], can achieve performance comparable to full-parameter fine-tuning by tuning less than 1% of the model parameters.
@@
In addition to the efficiency of fine-tuning, data availability has become another critical obstacle in the adaptation of PLM due to increasing considerations and regulations in data privacy [18, 19, 20]. Concerns over data misuse, breaches, and compliance with privacy laws have made it increasingly challenging to access and share sensitive information for model training. To address this, recent research has sought to integrate PEFT with Federated Learning (FL) [21, 22], which involves distributing the fine-tuning process across multiple decentralized devices or organizations, each training on their local data and sharing only the model update with a central server to create a global PEFT module. This synergy, often dubbed Federated PEFT (FedPEFT) [23], not only enables efficient adaptation of PLMs to specific tasks with low communication costs and computational overhead but also reduces the need for data centralization, mitigating privacy risks.
@@
Despite its merits, adapting PLMs in FL settings still faces security challenges that need to be addressed [23]. On the one hand, the PLM fine-tuning process introduces the risk of jailbreak, where malicious participants attempt to induce the models to generate harmful outputs in response to malicious prompts, against usage policies and social values [24]. Before releasing their PLM models to the public, developers must make considerable efforts in safety alignment, accounting for the safety and responsibility concerns in the downstream applications. For instance, Reinforcement Learning from Human Feedback (RLHF) [25] and RL from AI Feedback (RLAIF) [26] are well-established techniques that constrain the behaviors of PLMs according to human and AI preferences [27]. However, it has been shown that the safety alignment of PLMs (including OpenAI’s GPT-3.5 [3] and Meta’s LlaMA-2 [6]) can be compromised by fine-tuning with adversarially designed training examples [28, 29].
@@
This paper aims to raise awareness about the jailbreak threat associated with FedPEFT of PLMs, showing that PEFT methods can be exploited as an attack vector to circumvent the safety alignment of PLMs and generate harmful content against the usage policy and society in response to malicious prompts. We show that with less than 1% of the model’s parameters set as trainable, and a small subset of clients acting maliciously, the attack achieves surprisingly high attack success rates ($>80\%$)) with PEFT methods. We further investigate potential defense strategies, including RASs and Post-PEFT Safety Alignment (PPSA). The experimental results highlight the limitations of these defenses, particularly, even the advanced RASs, such as DnC [31], and ClippedClustering [30], struggle to defend against PaaA in scenarios with highly heterogeneous data distributions. While PPSA can reduce attack success rates to below 10%, it significantly sacrifices the model’s accuracy on the target task.
@@
These findings emphasize the urgent need for more advanced defense strategies that address the trade-offs between safety and utility in FedPEFT systems, offering insights for future research in this field. Moreover, our implementation has been integrated into Blades [^1] https://github.com/lishenghui/blades [34], a benchmark suite for FL security research, to facilitate reproducibility and stimulate further exploration within the community.
II Preliminaries
This section provides essential backgrounds on PLMs (Section II-A) and PEFT (Section II-B); the fundamentals of FL will be given in Section III.
II-A Pre-trained Language Models
PLMs have emerged as a cornerstone of modern NLP, demonstrating remarkable capabilities in understanding and generating human language [1, 2, 3, 4, 5, 6, 8, 9, 10]. These models leverage large-scale unsupervised pre-training on vast text corpora to learn rich contextual representations and linguistic patterns [35]. Through this process, PLMs acquire general language understanding that can be effectively transferred to various downstream tasks through fine-tuning or prompt-based approaches.
The architecture of contemporary PLMs predominantly follows the transformer-based design [36], which has proven highly effective in capturing long-range dependencies and contextual relationships in sequential data. The core building blocks of modern PLMs often consist of two fundamental components: Multi-Head Attention (MHA) and Feedforward Network (FFN), which are interconnected by residual connections and normalization layers [5, 6, 7]. MHA allows the model to simultaneously attend to different positions and aspects of the input sequence by projecting the input into multiple representation subspaces (heads), where each head can focus on different relationships within the data. FFN complements the attention mechanism with fully connected layers that introduce non-linear activation. Typically, it consists of two linear projection layers with an activation layer in between.
II-B Parameter-Efficient Fine-Tuning
(a) LoRA
(b) $(\text{IA})^{3}$
(c) LayerNorm
Figure 1: Architectures of the three PEFT methods examined in this paper. Trainable components are in orange, while frozen parameters are in light blue.
Aiming to achieve performance comparable to full-parameter fine-tuning while improving efficiency, PEFT methods propose to train only a small subset of the parameters while freezing the majority [13, 14]. In this paper, we focus on three representative PEFT methods listed below, including LoRA [14], $(\text{IA})^{3}$ [32], and LayerNorm [33]. Figure 1 illustrates the architectures of these methods.
LoRA [14]. This method reduces the number of trainable parameters by representing weight updates through low-rank decomposition across multiple model blocks and locations. Consider a PLM with $L$ LoRA modules, where weight matrices $\mathbf{W}_{i}\in\mathbb{R}^{m_{i}\times n_{i}}$ can be found in components such as MHA and FFN. For each weight matrix $\mathbf{W}_{i}$, LoRA introduces an update mechanism, $\mathbf{W}_{i}\leftarrow\mathbf{W}_{i}+\Delta\mathbf{W}_{i}$. The key idea is approximating the parameter update $\Delta\mathbf{W}_{i}$ through low-rank decomposition: $\Delta\mathbf{W}_{i}=\mathbf{A}_{i}\cdot\mathbf{B}_{i}^{\top}$, where $\mathbf{A}_{i}\in\mathbb{R}^{m_{i}\times k}$ and $\mathbf{B}_{i}\in\mathbb{R}^{n_{i}\times k}$ are trainable matrices, and $k\ll\min(m_{i},n_{i})$ is the reduced rank. This approach can be selectively applied to specific weight matrices in different model components, such as the query, key, and value projections in MHA, or the linear layers in FFN. Instead of updating the full weight matrix of size $m_{i}\times n_{i}$ (which requires $m_{i}\times n_{i}$ parameters), LoRA introduces only $k(m_{i}+n_{i})$ trainable parameters, enabling more efficient and targeted model adaptation.
$(\text{IA})^{3}$ [32]. This method modulates the internal activations of PLMs with lightweight scaling factors. For a given weight matrix $\mathbf{W}_{i}\in\mathbb{R}^{m_{i}\times n_{i}}$ and input $\mathbf{x}_{i}\in\mathbb{R}^{n_{i}}$, $(\text{IA})^{3}$ introduces a layer-specific scaling vector $\mathbf{l}_{i}\in\mathbb{R}^{n_{i}}$. The modified forward pass becomes:
$\rm{IA^{3}}(\mathbf{W}_{i},\mathbf{x}_{i})=\mathbf{l}_{i}\odot\gamma(\mathbf{W}_{i}\mathbf{x}_{i})$ (1)
where $\odot$ denotes element-wise multiplication, $\mathbf{l}_{i}$ acts as a learned scaling factor that can amplify or inhibit specific dimensions of the input activations, and $\gamma(\cdot)$ is defined as follows: when applied to FFNs, $\gamma$ is the activation function; when applied to MHAs, $\gamma$ is the identity function that outputs its input unchanged.
LayerNorm [33]. This method only fine-tunes the parameters of the normalization layers in a PLM to learn task-specific modifications. Specifically, the output of a normalization layer can be represented as:
$\rm{Norm}(\mathbf{x}_{i})=\frac{\mathbf{x}_{i}}{\sqrt{\frac{1}{d}\sum_{i=1}^{d}x_{i}^{2}}}\odot\mathbf{g}_{i},$ (2)
where $\mathbf{g}_{i}\in\mathbb{R}^{n_{i}}$ is a learnable scaling parameter, and $n_{i}$ denotes the dimensionality of the input. It is worth noting that such normalization differs from the classic LayerNorm [37]. Instead, we follow the RMSNorm [38], a variant that has become the de facto normalization in state-of-the-art PLMs [5, 6, 7, 8, 9, 10], including those examined in our study.
III System Model
@@
In this section, we introduce the system model studied in this work. Figure 2 shows an overview of the system model with numbered components (①–⑤) representing key elements.
III-A FedPEFT System
@@
$s_{k}^{i}=\rm{Template}(\text{cont}_{k}^{i},\text{inst}_{k}^{i},\text{resp}_{k}^{i}).$ (4)
@@ III-A1 FedPEFT Client
$\bm{\theta}_{k,t}^{s+1}={\rm{ClientOpt}}(\bm{\theta}_{k,t}^{s},\nabla\mathcal{L}_{k}(\bm{w},\bm{\theta}_{k,t}^{s}),\eta),$ (6)
@@
III-A2 FedPEFT Server
The server acts as the central coordinator in the training process. At the beginning of the training, it initializes the global PEFT parameters $\bm{\theta}_{0}$. For each communication round $t$, the server first selects a subset of clients $\mathcal{S}_{t}\subseteq[K]$ to participate in the current round. It then broadcasts the current global parameters $\bm{\theta}_{t}$ to these selected clients. Upon receiving the local updates from participating clients, the server aggregates them using the FedAvg algorithm [21] to obtain the new global parameters:
@@ III-A2 FedPEFT Server
III-B Threat Model
Our attack, dubbed “PEFT-as-an-Attack” (PaaA), aims to compromise the PLM’s safety guardrails through FedPEFT on malicious data. Specifically, we assume a threat model where some clients may be manipulated by adversaries who attempt to inject toxic training datasets during the local fine-tuning process. Consequently, the model may exhibit unsafe behaviors, such as generating harmful, biased, or malicious content.
Attack Objective. The primary objective of the adversaries is to bypass the pre-built safety guardrails of PLM and maximize the likelihood of harmful outputs conditioned on the corresponding harmful input instructions (as illustrated by ③ in Figure 2). Additionally, the adversaries strive to achieve this while maintaining stealth by ensuring that the global model’s performance on benign tasks remains unaffected, making the attack challenging to detect through standard validation datasets and metrics [40].
@@ III-C Defense Mechanisms
III-C1 Robust Aggregation Schemes
Traditionally, RASs (as illustrated by ④ in Figure 2) play an important role in mitigating the impact of adversarial updates by identifying and excluding their influence on the global model [43, 30]. Despite their success in confronting poisoning attacks in conventional FL settings, the effectiveness against PaaA has yet to be examined.
III-C2 Post-PEFT Safety Alignment
@@
While these conventional approaches were not specifically designed for our identified threat model, examining their effectiveness and limitations provides valuable insights into the challenges of defending against PaaA.
IV Experimental Setup
In this section, we present the experimental setup, including the system setup (Section IV-A), training details (Section IV-B), and the evaluation details (Section IV-C).
IV-A System Setup
@@
Table I: Overview of the selected PLMs and PEFT methods in our experiments. We also report trainable model parameters (# TPs), total parameters (# APs), and the ratio of trainable parameters.
@@
Figure 3: Performance comparison of three FedPEFT methods across 25 communication rounds, fine-tuned on the MedQA dataset without malicious clients. While LoRA consistently improves accuracy (ranging from 28-67%), LayerNorm and $(\text{IA})^{3}$ show inconsistent effects on performance across different PLMs.
IV-B Training Details
We adopt the Blades benchmark suite to simulate the FL system with adversarial settings [34]. For local training, we employ the SFTTrainer APIs provided by Huggingface PEFT [^2] https://huggingface.co/docs/peft/index and Unsloth [^3] https://unsloth.ai, and utilize the PLMs and datasets released on Huggingface. By leveraging the Bitsandbytes library [^4] https://huggingface.co/docs/bitsandbytes/main/en/index to perform 4-bit quantization of the PLMs, we are able to run all our experiments on a GPU cluster equipped with Nvidia Tesla A100 GPUs.
The AdamW [55] optimizer is applied with a batch size of 4. By default, the PEFT modules are trained for 25 communication rounds using FedAvg in our experiments (according to our evaluation, this number of rounds is sufficient to achieve performance improvement on the downstream tasks). We specify 50 local steps per round with the learning rate set to $0.001$.
IV-C Evaluation Details
@@
IV-C1 Safety
We generate responses for harmful instructions sourced from two representative safety benchmarks: AdvBench [57] and JailbreakBench [58]. Following prior art [59], we use OpenAI’s chat completion API [^5] https://openai.com/index/openai-api/ to judge whether each output is harmful or not and report the attack success rate (ASR), which is defined as the ratio of outputs that are judged as being harmful [^6] Note that the API is not perfect and provides only an approximate assessment of the potential harm..
IV-C2 Utility
We evaluate model utility on domain-specific downstream tasks using the test splits from MedQA and MetaMathQA, respectively. Both datasets come with unique correct answers, allowing for straightforward accuracy computation. We parse the model’s predicted answer for each response and compare it with the ground truth. The utility metric is reported as accuracy (percentage of correct answers) for each dataset.
V Experimental Results and Analysis
In this section, we begin by evaluating the effectiveness of three FedPEFT methods on MedQA without the presence of malicious clients (Section V-A). We then investigate the impact of PaaA on different FedPEFT methods (Section V-B). Finally, we examine the robustness of two sets of defense mechanisms, RASs (Section V-C) and PPSA (Section V-D), against PaaA.
V-A Learning Effectiveness of FedPEFT Methods
We first evaluate the effectiveness of different PEFT methods in adapting PLMs for MedQA under benign conditions (without malicious clients). Figure 3 shows the accuracy trajectories across 25 communication rounds on the MedQA dataset. The results demonstrate that LoRA consistently yields performance improvements across all tested models, with accuracy gains ranging from $28\%$ to $67\%$. In contrast, $(\text{IA})^{3}$ and LayerNorm show variable effects on model performance, sometimes even leading to degraded accuracy compared to the base models. While LoRA’s consistent improvements may stem from the low-rank decomposition, it is also important to note that LoRA typically involves more trainable parameters than the other methods, which may further contribute to its enhanced effectiveness.